When Raul and Rudy work together, they can paint a car in 8 days. Also, they could paint this car if Raul worked 12 days and Rudy worked 6 days. How long would it take each of them to paint the car alone?
Answer by stanbon(75887)

When Raul and Rudy work together, they can paint a car in 8 days. Also, they could paint this car if Raul worked 12 days and Rudy worked 6 days. How long would it take each of them to paint the car alone?
---
Let L and R be their respective rates
---
L + R = 1/8
12L + 6R = 1
-----------------------
Modify:
2L + 2R = 1/4
2L + R = 1/6
----------------
Subtract to solve for "R":
R = (6-4)/24 = 1/12
Rudy would take 12 days to do the job alone.
-----
Solve for "L":
L + R = 1/8
L = (1/8)-(1/12)
L = (12-8)/96
L = 4/96 = 1/24
Raul would take 24 days to do the job alone.
------
Cheers,
Stan H.